OpenStudy (anonymous):

solve 3a(3a-x)=2b(x+2b)

OpenStudy (shamim):

\[3a(3a-x)=2b(x+2b)\]

OpenStudy (shamim):

\[or, 9a ^{2}-3ax=2bx+4b ^{2}\]

OpenStudy (shamim):

\[3ax+2bx=9a ^{2}-4b ^{2}\]

OpenStudy (shamim):

\[x(3a+2b)=(3a)^{2}-(2b)^{2}\]

OpenStudy (shamim):

\[x(3a+2b)=(3a+2b)(3a-2b)\]

OpenStudy (shamim):

\[or, x=3a-2b\]

OpenStudy (anonymous):

thanks. for helping out

OpenStudy (shamim):

u r most welcome